# itemizestditemize

### v1

@@ V-B A Priori Suboptimality Bound of RAG
We present the a priori suboptimality bound of RAG. The bound has theoretical and practical value. On the theoretical side, it captures the suboptimality cost due to decentralization as a function of the overlap of each robot $i$’s action with the actions of its non-neighbors, i.e., the robots that robot $i$ does not coordinate with, as captured by ${\sf coin}_{f,i}$. On the practical side, by bounding ${\sf coin}_{f}$ with a computable bound as a function of the agents’ neighborhoods, we enable the agents to optimize their neighborhoods to maximize the suboptimality bound of RAG subject to their communication-bandwidth constraints. An example of such an upper bound was given in Example 1.
@@
$f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,\frac{1}{2}\left[f(\,\mathcal{A}^{\sf OPT}\,)-\sum_{i\,\in\,{\cal N}}{\sf coin}_{f,i}(\,{\cal N}_{i}\,)\right].$ (9)
Theorem 1 captures the intuition that when the agents coordinate with fewer other agents (more decentralization), then the approximation performance will deteriorate. This intuition is made rigorous by applying Proposition 1 in eq. 9 along the spectrum of increasing decentralization from fully centralized to fully decentralized networks:
@@
    $f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,\frac{1}{2}\,f(\,\mathcal{A}^{\sf OPT}\,),$ (10)
    i.e.RAG is near-optimal, matching the approximation ratio $1/2$ of the seminal SG algorithm [18]. The bound $1/2$ is near-optimal since the best possible bound for the optimization problem in (3) is $1-1/e\simeq 0.63$ [16].
    Equation 10 is derived from eq. 9 since for fully centralized networks ${\cal N}_{i}={\cal N}\setminus\{i\}$ and ${\sf coin}_{f,i}({\cal N}\setminus\{i\})=0$, for all $i\in{\cal N}$ (Proposition 1).
    Remark 4 (Decision time).
    For full centralization, RAG will be the slowest among all possible communication-network topologies, with an $O(|{\cal N}|)$ decision time. The reason is that the agents will be choosing actions sequentially (no parallelization is possible). Still, RAG will be as fast or faster than the state of the art, as we elaborate on in Section VI.
@@
    Remark 5 (Decision time).
    Across any network topology, RAG runs up to two orders faster than the state of the art, as we elaborate in Section VI and summarize in Table I.
@@
    $f(\,{\cal A}^{{\sf RAG}}\,)$ $\,\geq\,\frac{1}{2}\left[f(\,\mathcal{A}^{\sf OPT}\,)-\sum_{i\,\in\,{\cal N}}{\sf coin}_{f,i}(\,\emptyset\,)\right]$ (11)
    $\,\in\,\left[\frac{1-\kappa_{f}}{2-\kappa_{f}},\,\frac{1}{2}\right]f(\,\mathcal{A}^{\sf OPT}\,).$ (12)
    Equation 11 captures the intuition that when the agents’ actions do not overlap, then no communication still leads to near-optimal performance. For example, per the area coverage Example 1, when the agents are sufficiently far away such that their field of views cannot overlap upon executing their actions, then ${\sf coin}_{f,i}(\,\emptyset\,)=0$ for all $i\in{\cal N}$. Particularly, then the bound in eq. 11 becomes $1/2$, matching the fully centralized performance.
    In the worst case, the bound in eq. 9 takes the value $(1-\kappa_{f})/(2-\kappa_{f})$, and becomes zero when the actions of all agents fully overlap with each other ($\kappa_{f}=1$). This is inevitable since all agents ignore all others and thus cannot coordinate actions to reduce the overlap.
    Equation 11 is derived from eq. 9 since for fully decentralization networks ${\cal N}_{i}=\emptyset$, for all $i\in{\cal N}$. Equation 12 is derived from eq. 11 by substituting the worst-case bound for ${\sf coin}_{f,i}(\emptyset)$, that is, ${\sf coin}_{f,i}(\emptyset)\leq\kappa_{f}\,f(a_{i})$ (Proposition 1).
    Remark 6 (Decision time).
    RAG runs in $O(\tau_{f}|{\cal V}_{i}|)$ time in this case since the agents simply need to choose their best available action without coordinating with any agent. RAG takes the same decision time as the state of the art since all methods just require one time of myopic action selection.
The tightness of the bound in Theorem 1 will be analyzed in our future work.
@@ V-C A Posteriori Suboptimality Bound of RAG
We present the a posteriori approximation bound of RAG, i.e., the bound that is computable after all robots have selected actions (Theorem 2). We also prove that the bound is non-decreasing and submodular as a function of the agents’ neighborhoods (Proposition 2).
The practical implications of Proposition 2 are dual: First, relatively smaller neighborhoods can marginally achieve the same performance as large neighborhoods; equivalently, decentralized coordination can marginally be as near-optimal as centralized coordination (Remark 7). Second, larger neighborhoods do not necessarily result in better coordination performance when we account for both the decision time and the suboptimality of the algorithm (Remark 8). We validate the above in the experimental results of Section VII.
To state the first theorem, we recall the notation:
@@
Given the actions $\{a_{i}^{\sf RAG}\}_{i\,\in\,{\cal N}}$ selected by the agents, RAG guarantees [^2] Theorem 2 holds true for $f$ non-decreasing and submodular and not necessarily 2nd-order submodular.
$\hskip 8.53581ptf(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,f(\,\mathcal{A}^{\sf OPT}\,)\,-\,\sum_{i\,\in\,{\cal N}}f(\,a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal I}_{i}}^{\sf RAG}\,).$ (13)
Theorem 2 captures the suboptimality cost due to decentralization, similarly to Theorem 1. But in contrast to Theorem 1, Theorem 2 captures the decentralization cost as a function of the action overlap between agent $i$ and its neighbors, instead of its non-neighbors. As such, Theorem 2 captures the intuition that the larger agent $i$’s neighborhood is, then the better the suboptimality guarantee can be since then agent $i$ would have the chance to coordinate actions with more agents. This intuition is made rigorous with the following theorem.
@@
$f(\,\mathcal{A}^{\sf OPT}\,)\,-\,\sum_{i\,\in\,{\cal N}}f(\,a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal I}_{i}}^{\sf RAG}\,)$ is non-decrea- sing and approximate submodular as a function of $\{{\cal I}_{i}\}_{i\,\in\,{\cal N}}$.
The proposition implies that although the approximation performance will indeed improve if the agents have more neighbors, the marginal improvement diminishes. This observation has the interwoven practical implications remarked below.
Remark 7 (Smaller Neighborhoods Can Be Enough).
Since making the neighborhoods larger has a diminishing returns effect in the suboptimality performance, beyond a large enough size of neighborhood, there is not much more to gain. Therefore, decentralized coordination can marginally be as near-optimal as centralized coordination.
Remark 8 (Larger Neighborhoods Disproportionally Impact Decision Time).
When we account for the impact that the size of the neighborhoods can have to the trade-off of decision time with optimality, larger neighborhoods are not necessarily better. Intuitively, larger neighborhoods will lead to better performance in terms of maximizing the global objective $f$. But this may come at the cost of disproportionally more time spent for coordination. The reason is that the gain in the suboptimality guarantee diminishes as the neighborhoods become larger, per Proposition 2, i.e., it is sublinear with the size of the neighborhoods. But the decision time may increase linearly for RAG and superlinearly for the state of the art (Section VI). Therefore, the benefit of the marginal increase in the suboptimality guarantee may be negatively outweighed by a large increase in the decision time of the algorithm.
We validate both remarks in the experimental results of Section VII over a scenario of road coverage (Fig. 8).
@@ V-D Comparison to the State of the Art
RAG $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1/2$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1/2$               $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1/2$
Figure 6: Comparison of Suboptimality Guarantees of [31, 32] and RAG: Example of Area Coverage. We make the comparison over an area coverage task with multiple drones, as in Fig. 5(a), where the drones need to each select a short trajectory to maximize the total covered area at the next time step. The stars are the drones’ positions, the circles are the their field-of-views (FOVs), and the red lines denote undirected communication links. The best performances for each metric are colored in blue. (a) The drones are fully decentralized (no inter-agent communication). The bounds by [31, 32] are both $1/4$. For $|{\cal N}|$ agents, instead of $4$, the bounds would be $O(1/|{\cal N}|)$. However, since the drones are far enough for their FOV to overlap, ${\sf coin}_{f,i}=0,\forall i$, and thus RAG can guarantee $1/2$. (b) The drones are partially decentralized such that only the physically close drones communicate. For this communication graph, which we assume to be the same as the information-action access graph, $\omega=2$, $\chi=2$, and $\alpha^{\star}=3$. Therefore, [31] gives $[1/4,1/2]$ and [32] gives $[1/4,1/3]$. RAG still gives $1/2$ since non-neighbors are far away for all drones. (c) In the fully centralized case, all methods give $1/2$. In all, RAG may provide tighter bounds than [31, 32] since it considers the actual function $f$ using ${\sf coin}_{f}$ but [31, 32] are agnostic to $f$.
We summarize the approximation guarantees of the state of the art and of RAG in Table I. Therein, we observe the trade-off of decision time and optimality: the algorithms with the best suboptimality guarantees —first six rows of the table, achieving the near-optimal $1/2$ or $1-1/e$ [16]— can exhibit also the worst decision times, one to two orders higher than that of RAG. We discuss the algorithms’ decision times in more detail in Section VI. Among the remaining algorithms —three last rows of the table— RAG is the only algorithm that provides task-aware ($f$-based) performance guarantees, and that quantifies the suboptimality guarantee as a function of each agent’s local communication network of agents. Instead, the guarantees of [31, 32] are task-agnostic and can scale inversely proportional to the number of agents even when RAG can still guarantee the near-optimal $1/2$, as illustrated in Fig. 6.
@@
where ${\cal N}_{i}^{\text{in}}\subseteq[i-1]$. Equation 14 generalizes SG’s rule in eq. 4 to the setting where agent $i$ has access only to the actions selected by the agents in ${\cal N}_{i}^{\text{in}}\subseteq[i-1]$, instead of all agents that have selected an action before agent $i$. The information-access structure prescribed by the rule in eq. 14 can be represented as a directed acyclic graph ${\cal G}_{\text{info}}$, where agent $i$’s neighbors in ${\cal G}_{\text{info}}$ are the set ${\cal N}_{i}^{\text{in}}$ of agents. [^3] The graph ${\cal G}_{\text{info}}$ is, in general, different from the communication graph ${\cal G}$. That is, when an agent $i$ and an agent $j$ do not communicate (they are not neighbors in the communication graph ${\cal G}$) but agent $j\in{\cal N}_{i}^{\text{in}}$, then, agent $j$’s action needs to be relayed to agent $i$ via other agents in ${\cal G}$ that form a connected communication path in ${\cal G}$ between agent $i$ and agent $j$. Due to the limited information access, the suboptimality guarantees of [31, 32] take the form presented in Table I, where $\omega({\cal G}_{\text{info}})$ is the clique number of ${\cal G}_{\text{info}}$, $\chi({\cal G}_{\text{info}})$ is the chromatic number, and $\alpha^{\star}({\cal G}_{\text{info}})$ is the fractional independence number [49].
@@ VI Decision Time Analysis
We present the decision time of RAG, that is, the time it takes for RAG to terminate, and provide an approximate solution to 1. RAG’s decision time scales linearly with the size of the network, up to two orders faster than the state-of-the-art algorithms. We summarize the decision time of the state of the art and of RAG in Table I, where we use the notation:
@@ VI-A Decision Time of RAG
We next first analyze the computational and communication complexities of RAG and then present its decision time.
@@ VI-B Comparison to the State of the Art
We summarize the decision times of the state of the art and of RAG in Table I. RAG’s decision time scales linearly with the network’s size, namely, $|{\cal N}|$, whereas, in the worst case, the state of the art scales at least quadratically with $|{\cal N}|$. Specifically, RAG has computational time that is linear in $|{\cal N}_{i}|$, independent of $|{\cal N}|$, and communication time linear in $|{\cal N}|$. The comparison is summarized in Table I. Therein, for the sake of the comparison, we assume for simplicity that $|{\cal V}_{i}|\,=\!|{\cal V}_{j}|,\forall i,j\in{\cal N}$. Also, we divide the state of the art into algorithms that solve 1 either indirectly in the continuous domain via employing the multi-linear extension [48] of the set function $f$ [20, 47, 46], or directly in the discrete domain [39, 21, 43, 31, 32]: [^4] The continuous-domain algorithms employ consensus-based techniques [20, 46], or algorithmic game theory [47], and need to compute the multi-linear extension’s gradients via sampling., [^5] The decision times of the continuous-domain algorithms depend on additional problem-dependent parameters (such as Lipschitz constants, the diameter of the domain set of the multi-linear extension, and a bound on the gradient of the multi-linear extension), which we make implicit in Table I via the $O$, $\Omega$, and $\Theta$ notations., [^6] The computational and communication times reported for [47] are based on the numerical evaluations therein since a theoretical quantification is missing in [47] and appears non-trivial to derive one as a function of ${\cal N}$, $\epsilon$, or any other of the problem parameters.
@@
RAG requires at most $(\tau_{c}+\tau_{\#})\,(|{\cal N}|-1)$ communication time. In [31, 32], the agents need to communicate over ${\cal G}$ per the pre-specified information-access directed acyclic graph ${\cal G}_{\text{info}}$ per the rule in eq. 14 [31, Remark 3.2]. In Appendix IV, we identify a worst case where this rule results in $O(\tau_{c}\,|{\cal N}|^{2})$ communication time. This happens when each agent $i-1$ and agent $i$ in the decision sequence do not communicate directly, thus, the information of agent $i-1$ needs to be relayed to agent $i$ via other agents that form a connected communication path between the two. The method in [21], which introduces a depth-first search (DFS) procedure to determine the best agents’ ordering to run SG [18] over arbitrary (strongly) connected networks (instead of just line graphs), requires a worst-case communication time of $O(\tau_{c}\,|{\cal N}|^{3})$ for directed networks, and $O(\tau_{c}\,|{\cal N}|^{2})$ for undirected networks [6, Appendix II]. For some $\epsilon$, the method in [39] may require less communication time for running the algorithm per se than other methods, but an additional communication time of $O(\tau_{c}\,|{\cal V}_{i}|)$ is needed to distributively find a DAG that optimizes the algorithm’s approximation performance. The remaining algorithms require communication times proportional to $|{\cal N}|^{2}$ or more.

### v2

@@ V-B A Priori Suboptimality Bound of RAG
We present the a priori suboptimality bound of RAG. by bounding ${\sf coin}_{f}$ with a computable bound as a function of the agents’ neighborhoods, we enable the agents to optimize their neighborhoods to maximize the suboptimality bound of RAG subject to their communication-bandwidth constraints.
@@
$f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,\frac{1}{1+\kappa_{f}}\left[f(\,\mathcal{A}^{\sf OPT}\,)-\kappa_{f}\sum_{i\in{\cal N}}{\sf coin}_{f,i}(\,{\cal N}_{i}\,)\right].$ (12)
Theorem 1 captures the intuition that when the agents coordinate with fewer other agents, then the approximation performance will deteriorate. This intuition is made rigorous by applying Proposition 1 to eq. 12 along the spectrum from fully centralized to fully decentralized networks:
@@
    $f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,\frac{1}{1+\kappa_{f}}\,f(\,\mathcal{A}^{\sf OPT}\,),$ (13)
    i.e.RAG is near-optimal, matching the approximation ratio $1/(1+\kappa_{f})$ of the seminal SG algorithm [50]. The bound $1/(1+\kappa_{f})$ is near-optimal since the best possible bound for the optimization problem in (6) is $1-\kappa_{f}/e$ [10].
@@
    $f(\,{\cal A}^{{\sf RAG}}\,)$ $\,\geq\,\frac{1}{1+\kappa_{f}}\left[f(\,\mathcal{A}^{\sf OPT}\,)-\kappa_{f}\sum_{i\in{\cal N}}{\sf coin}_{f,i}(\,\emptyset\,)\right]$ (14)
    $\,\in\,\left[1-\kappa_{f},\,\frac{1}{1+\kappa_{f}}\right]f(\,\mathcal{A}^{\sf OPT}\,).$ (15)
    Equation 14 captures the intuition that when the agents’ actions do not overlap, then no communication still leads to near-optimal performance. For example, per the area coverage Example 1, when the agents are sufficiently far away such that their field of views cannot overlap upon executing their actions, then ${\sf coin}_{f,i}(\,\emptyset\,)=0$ for all $i\in{\cal N}$. Particularly, then the bound in eq. 14 becomes $1/(1+\kappa_{f})$, matching the fully centralized performance.
    In the worst case, the bound in eq. 12 takes the value $1-\kappa_{f}$, and becomes zero when the actions of all agents fully overlap with each other ($\kappa_{f}=1$). This is inevitable since all agents ignore all others and thus cannot coordinate actions to reduce the overlap.
The tightness of the bound will be analyzed in future work.
Corollary 1 (A Priori Bound with Approximate Greedy Selection).
If Algorithm 1’s line 3 can only perform approximate greedy selection such that $f(a_{i})\geq\eta\,f(a_{i}^{\sf OPT}\,|\,{\cal A}_{i}),0<\eta\leq 1$, then RAG guarantees:
$f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,$ $\frac{\eta}{1+\eta\kappa_{f}}\Bigg{[}f(\,\mathcal{A}^{\sf OPT}\,)$
$-\left(\frac{1}{\eta}-1+\kappa_{f}\right)\sum_{i\in{\cal N}}{\sf coin}_{f,i}(\,{\cal N}_{i}\,)\Bigg{]}.$ (16)
Also, it holds that:
$\frac{f(\,{\cal A}^{{\sf RAG}}\,)}{f(\,{\cal A}^{\sf OPT}\,)}\geq\begin{cases}\frac{\eta}{1+\eta\kappa_{f}},&{\cal G}\text{ is {fully centralized}},\\ \eta(1-\kappa_{f}),&{\cal G}\text{ is {not fully centralized}}.\end{cases}$ (17)
RAG still has near-optimal performance guarantees even with approximate greedy selection [13, 15]. When $\eta=1$, Corollary 1 reduces to Theorem 1.
@@ V-C A Posteriori Suboptimality Bound of RAG
We present the a posteriori approximation bound of RAG (Theorem 2). We recall the notation:
@@
Given the actions $\{a_{i}^{\sf RAG}\}_{i\,\in\,{\cal N}}$ selected by the agents, RAG guarantees: [^3] Theorem 2 holds true for $f$ non-decreasing and submodular and not necessarily 2nd-order submodular.
$\hskip 8.53581ptf(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,f(\,\mathcal{A}^{\sf OPT}\,)\,-\,\kappa_{f}\sum_{i\in{\cal N}}f(\,a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal I}_{i}}^{\sf RAG}\,).$ (18)
Theorem 2 captures the suboptimality cost due to decentralization, similarly to Theorem 1. In contrast to Theorem 1, Theorem 2 captures the decentralization cost as a function of the action overlap between agent $i$ and its neighbors, instead of its non-neighbors. As such, Theorem 2 captures the intuition that the larger agent $i$’s neighborhood is, the better the suboptimality guarantee can be since then agent $i$ would have the chance to coordinate actions with more agents.
@@
The right-hand side of eq. 18 is non-decreasing and approximate submodular as a function of $\{{\cal I}_{i}\}_{i\,\in\,{\cal N}}$.
The proposition implies that although the approximation performance will improve if the agents have more neighbors, the gained improvement diminishes.
Remark 5 (Trade-Off between Decision Speed and Optimality).
For larger neighborhoods, the marginal increase in the approximation guarantee is negatively outweighed by a greater increase in decision time. The reason is that the gain in the approximation guarantee diminishes as the neighborhoods become larger (Proposition 2) while the decision time may increase linearly for RAG and superlinearly for the state of the art, as we present in Section VI.
The appropriate size of neighborhoods that balances the trade-off for the task at hand can be found via experimentation in high-fidelity simulations, as we illustrate in the experiments.
@@ V-D Comparison to the State of the Art
[…]
Figure 6: Comparison of Suboptimality Guarantees of [37, 38] and RAG: Example of Area Coverage. We make the comparison over an area coverage task with multiple drones, as in Fig. 5(a), where the drones need to each select a short trajectory to maximize the total covered area at the next time step. The stars are the drones’ positions, the circles are their field-of-views (FOVs), and the red lines denote undirected communication links. The best performances for each metric are colored in blue. (a) The drones are fully decentralized (no communication). The bounds by [37, 38] are both $1/4$. For $|{\cal N}|$ drones, instead of $4$, the bounds would be $O(1/|{\cal N}|)$. Instead, RAG guarantees 1/($11/(1+\kappa_{f})=$ 1 since the drones are far enough for their FOV to overlap, ${\sf coin}_{f,i}=0,\forall i$, and $\kappa_{f}=0$. (b) The drones are partially decentralized such that only the physically close drones communicate. For this communication graph, that we assume to be the same as the information-action access graph, $\omega=2$, $\chi=2$, and $\alpha^{\star}=3$. Therefore, [37] gives $[1/4,1/2]$ and [38] gives $[1/4,1/3]$. RAG still gives $21/(1+\kappa_{f})\geq 1/$ 2 since non-neighbors are far away. (c) In the fully centralized case, [37, 38] give $1/2$ and RAG gives $1/(1+\kappa_{f})$. In all, RAG may provide tighter bounds than [37, 38] since it considers the actual function $f$ using ${\sf coin}_{f}$ and $\kappa_{f}$, whereas [37, 38] are agnostic to $f$.
We summarize the approximation guarantees of the state of the art and RAG in Table I. We observe the trade-off of decision time and optimality: the algorithms with the best suboptimality guarantees —first six rows of the table, achieving the near-optimal $1/2$ or $1-1/e$ [10]— can exhibit also the worst decision times, one to two orders of the network size higher than that of RAG. Among the remaining algorithms —three last rows of the table— RAG is the only algorithm that provides task-aware ($f$-based) performance guarantees and that quantifies the suboptimality guarantee as a function of each agent’s local communication network. Instead, the guarantees of [37, 38] are task-agnostic and can scale inversely proportional to the number of agents even when RAG can still guarantee the near-optimal $1/2$ (Fig. 6).
@@
where ${\cal N}_{i}^{\text{in}}\subseteq[i-1]$. Equation 19 generalizes SG’s rule in eq. 7 to the setting where agent $i$ has access only to the actions selected by the agents in ${\cal N}_{i}^{\text{in}}\subseteq[i-1]$, instead of all agents that have selected an action before agent $i$. The information-access structure prescribed by the rule in eq. 19 can be represented as a directed acyclic graph ${\cal G}_{\text{info}}$, where agent $i$’s neighbors in ${\cal G}_{\text{info}}$ are the set ${\cal N}_{i}^{\text{in}}$ of agents. [^4] The graph ${\cal G}_{\text{info}}$ is in general different from the communication graph ${\cal G}$. When an agent $i$ and an agent $j$ do not communicate (they are not neighbors in the communication graph ${\cal G}$) but agent $j\in{\cal N}_{i}^{\text{in}}$, then agent $j$’s action needs to be relayed to agent $i$ via other agents in ${\cal G}$ that form a connected communication path in ${\cal G}$ between agent $i$ and agent $j$. Due to the limited information access, the suboptimality guarantees of [37, 38] take the form presented in Table I, where $\omega({\cal G}_{\text{info}})$ is the clique number of ${\cal G}_{\text{info}}$, $\chi({\cal G}_{\text{info}})$ is the chromatic number, and $\alpha^{\star}({\cal G}_{\text{info}})$ is the fractional independence number [55].
@@ VI Decision Time Analysis
We bound the time it takes for RAG to terminate. RAG’s decision time scales linearly with the size of the network, up to two orders of the network size faster than the state of the art. We summarize the decision time of the state of the art and of RAG in Table I, where we use the notation:
@@ VI-A Decision Time of RAG
We first analyze the computational and communication complexities of RAG and then present its decision time.
@@ VI-B Comparison to the State of the Art
We summarize the decision times of the state of the art and RAG in Table I. RAG scales linearly with the number of robots, $|{\cal N}|$, whereas the state of the art scales at least quadratically with $|{\cal N}|$. RAG has computational time that is linear in $|{\cal N}_{i}|$, independent of $|{\cal N}|$, and communication time linear in $|{\cal N}|$.
In Table I, we assume for simplicity that $|{\cal V}_{i}|\,=\!|{\cal V}_{j}|,\forall i,j\in{\cal N}$. We divide the state of the art into algorithms that solve 1 either in the continuous domain via employing the multi-linear extension [54] of the set function $f$ [30, 53, 52], or in the discrete domain [44, 31, 48, 37, 38]: [^5] The continuous-domain algorithms employ consensus-based techniques [30, 52], or algorithmic game theory [53], and need to compute the multi-linear extension’s gradients via sampling., [^6] The decision times of the continuous-domain algorithms depend on additional problem-dependent parameters (such as Lipschitz constants, the diameter of the domain set of the multi-linear extension, and a bound on the gradient of the multi-linear extension), which we make implicit in Table I via the $O$, $\Omega$, and $\Theta$ notations., [^7] The computational and communication times reported for [53] are based on the numerical evaluations therein since a theoretical quantification is not included in [53] and appears non-trivial to derive one as a function of ${\cal N}$, $\epsilon$, or the other problem parameters.
@@
RAG requires at most $(\tau_{c}+\tau_{\#})\,(|{\cal N}|-1)$ communication time. In [37, 38], the agents need to communicate over ${\cal G}$ per the pre-specified information-access directed acyclic graph ${\cal G}_{\text{info}}$ per the rule in eq. 19 [37, Remark 3.2]. In Appendix IV, we identify a worst case where this rule results in $O(\tau_{c}\,|{\cal N}|^{2})$ communication time. This happens when each agent $i-1$ and agent $i$ in the decision sequence do not communicate directly, thus, the information of agent $i-1$ needs to be relayed to agent $i$ via other agents that form a connected communication path between the two. The method in [31], which introduces a depth-first search (DFS) procedure to determine the best agents’ ordering to run SG [11] over arbitrary (strongly) connected networks (instead of just line graphs), requires a worst-case communication time of $O(\tau_{c}\,|{\cal N}|^{3})$ for directed networks, and $O(\tau_{c}\,|{\cal N}|^{2})$ for undirected networks [28, Appendix II]. For some $\epsilon$, the method in [44] may require less communication time for running the algorithm per se than other methods, but an additional communication time of $O(\tau_{c}\,|{\cal V}_{i}|)$ is needed to distributively find a DAG that optimizes the algorithm’s approximation performance. The remaining algorithms require communication times proportional to $|{\cal N}|^{2}$ or more.